# Learning to chain-of-thought with Jensen’s evidence lower bound

### v1

@@ 4 Tightening the objective via multi-sample Jensen’s lower bound
5:     (iii) Update $\theta$ with $n$-sample average of gradient estimate Eqn (5) or its multi-sample variant Eqn (7).
@@
Algorithm 1 Chain-of-thought optimization with Jensen’s lower bound (or its multi-sample extension)
5 Connections to algorithmic alternatives
The lower bound approach bears close connections to a number of algorithmic alternatives, which we discuss below. See Algorithm 1 for the pseudocode for of full algorithm.
5.1 Evidence lower bound
The evidence lower bounds (ELBO) (Blei and Jordan, 2006; Kingma and Welling, 2013a; Burda et al., 2015) control for the tightness of the lower bound with an inference distribution $q_{\phi}(c|x,a^{\ast})$ which defines a distribution over chain-of-thoughts. The lower bound takes the following form
$\mathcal{L}_{\theta,\phi}(x,a^{\ast})=\mathbb{E}_{c}\left[\log\pi_{\theta}(a^{\ast}|x,c)-\log\frac{q_{\phi}(c|x,a^{*})}{\pi_{\theta}(c|x)}\right],$
where the expectation is under $c\sim q_{\phi}(\cdot|x,a^{\ast})$. It lower bounds the marginal log likelihood $\mathcal{L}_{\theta,\phi}(x,a^{\ast})\leq\log\pi_{\theta}(a^{\ast}|x)$ and it is tight if and only if the inference distribution equals the posterior distribution $q_{\phi}(c|x,a^{\ast})=p^{\pi_{\theta}}(c|x,a^{\ast})$. Since ELBO is a function of both policy parameter $\theta$ and inference distribution parameter $\phi$, we can optimize both with stochastic gradient estimates: given a chain-of-thought sample $c\sim q_{\phi}(\cdot|x,a^{\ast})$,
$\ \ -\nabla_{\phi}\log q_{\phi}(c|x,a^{*}).$
Juxtaposing the form of the gradient here and the gradient to the Jensen’s lower bound defined in Eqn (5), we observe that the inference distribution gradient $g_{\phi}$ bears resemblance to the REINFORCE gradient; while the policy distribution gradient $g_{\theta}$ bears resemblance to the SFT gradient. In fact, we can show that under the special parameterization $q_{\phi}(c|x,a^{\ast})\coloneqq\pi_{\theta}(c|x)$, the two gradients are exactly equivalent. Such an observation is stated formally below.
Lemma 1.
(Jensen’s lower bound as a special case of ELBO) When $q_{\phi}(c|x,a^{\ast})\coloneqq\pi_{\theta}(c|x)$, ELBO is equivalent to the Jensen’s lower bound $\mathcal{L}_{\theta,\phi}(x,a^{\ast})=\mathcal{L}_{\theta}(x,a^{\ast})$ gradient esimtates are equivalent to the Jensen’s lower bound’s stochastic gradient estimates.
Proof.
The proof follows from the fact that when $q_{\phi}=\pi_{\theta}$, we have
By introducing an inference distribution $q_{\phi}$, ELBO is much expressive than the Jensen’s lower bound and allows for a tighter approximation to the marginal log likelihood. However, this also introduces additional complexity of having to learn the approximate posterior distribution. In our applications of interest, training a posterior model of a large size can be a major computational overhead. In practice, Hoffman et al. (2024) approximates the posterior via a few steps of MCMC and forgoes learning with such a distribution altogether. We take a different approach with a similar motivation: by tightening the lower bound with multiple samples, we also avoid the need for an explicit posterior.
5.2 Reinforcement learning
We show that there is a close connection between the lower bound formulation and the expected reward (return) maximization objective in RL (Sutton and Barto, 1998). Concretely, we will see how the lower bound objectives are closely related to a conditional expectation trick that produces a RL policy gradient estimate with lower variance. First, we show that (up to a log transform) RL optimizes for the same target as the lower bound objectives, given the indicator reward.
Lemma 2.
Proof.
Assuming access to $n$ i.i.d. trajectories $(y_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)$, we start with the classic RL policy gradient with leave-one-out baseline (i.e., RLOO (Ahmadian et al., 2024b))
@@ 5.2 Reinforcement learning
The proof is provided in Appendix C. Putting $g_{\text{var-reduced-pg}}$ from Eqn (9) and the gradient estimate of the Jensen’s lower bound (Eqn (5)) side-by-side, we identify intriguing similarities. Both gradient estimates employ two terms that update either the chain-of-thought component $\pi_{\theta}(\cdot|x)$ or the answer component $\pi_{\theta}(\cdot|x,c)$, with the only subtle difference being the extra log-transform needed for obtaining the Jensen lower bound. This alludes to the fact that the lower bound gradient has intrinsic built-in variance reduction.
We provide additional discussions of a few practical trade-offs in using the variance-reduced estimate $g_{\text{var-reduced-pg}}$ in Appendix C.
Figure 2: Short-form answer experiments with MATH. We compare three baselines: online RL with access to the oracle Sympy-based reward and lower bound algorithms. In the left plot, we monitor the reward on the training dataset. Online RL obtains the best training time trade-off, followed by multi-sample lower bound and the single-sample lower bound; In the middle plot, we monitor the evaluation on a test set during training. Multi-sample lower bound and online RL obtains similar performance; In the right plot, we graph training reward against the lower bound objectives, averaged over training tokens. The two signals bear positive correlations overall and multi-sample lower bound yields better correlations.
@@ 5.3 Optimizing Jensen’s lower bound with regularization is optimizing a special ELBO
Applying the same derivation as the regular ELBO, log likelihood $\log p_{\pi_{\theta},\pi_{\text{ref}}}(a|x)$ is lower bounded as
@@
Note that the aggregate objective Eqn (12) can also be optimized via stochastic gradient ascent. We just need to add an additional term associated with the KL regularization, to the original gradient estimate to $\mathcal{L}_{\theta}(x,a^{*})$ defined in Eqn (5). An example of such a gradient estimate usually takes the following form
@@
6 Lower bound vs. RL: practical trade-offs
In this section, we will discuss a few practical trade-offs of different algorithms.
6.1 Choice of reward
@@
A key difference between RL and lower bound methods is the choice of reward. As discussed earlier, lower bound methods do not require a reward - more concretely, they can be understood as adopting the indicator reward $r(x,y)=\mathds{1}_{\{a=a^{\ast}\}}$. In practice, this can be instantiated as a strict string match float(answer == gt_answer). However, such a reward function will likely induce false negatives, as semantically equivalent generations might be vastly different strings. In practice, a more lenient match is typically applied to better balance the false negative. For example, for math problems (Hendrycks et al., 2021; Yue et al., 2024), usually programmtic checks are implemented to check for equivalence of two short expressions, such that e.g., pi and 3.1415926 might be considered equivalent.
@@
More formally, consider a general reward function $r(x,y)=\text{match}(a,a^{*})$ calculated as a binary match between $a$ and $a^{*}$. We can rewrite the RL objective as $\mathbb{E}[\text{match}(a,a^{*})]$. In order to adapt the formulation in this work to the lower bound case, we need to work explicitly with the equivalent set $\mathcal{A}\coloneqq\{a|\text{match}(a,a^{*})=1\}$. We will need to calculate quantities such as the probability $\pi_{\theta}(\mathcal{A}|x,c)\coloneqq\sum_{a\in\mathcal{A}}\pi_{\theta}(a|x,c)$, which reduces to $\pi_{\theta}(a^{*}|x,c)$ in case we use exact match. Computing such probabilities is expensive since we need to enumerate all $a\in\mathcal{A}$ if inverting the match function is feasible at all. As a result, the lower bound formulation cannot be adapted to generic match function or reward function.
In summary, when a good reward is available, online RL is at an advantage due to its flexibility. There are also cases where good rewards are not easy to come by, and adopting the lower bound formulation is a decent default strategy. An example is where the ground truth answer takes a rather long form, in which case string match or programmatic check will produce too much false negatives.
6.2 Further details of lower bound algorithms
One technical challenge of the lower bound algorithm is that, given a full generation $y$, we need to define the boundary between the chain-of-thought $c$ and answer $a$ such that $y=[c,a]$ where we can interpret the latter as the concatenation of two arrays of tokens (or strings). Usually, we can ask the model to generate responses in a particular format, e.g., concluding the generation with a phrase the final answer is (Yue et al., 2024). We can use this phrase as an identifier, and define everything after the phrase as the answer $a$ and otherwise as the chain-of-thought (this step is part of the algorithm pseudocode in Algorithm 1). During training, we will need to enforce such formatting from the model so that the loss can be applied properly.
One mitigating strategy is to apply the KL regularization against the reference policy. The initial motive for such regularization was to prevent the model from degrading its general capability during RL finetuning (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). We will ablate how useful such regularization is to the lower bound methods. Besides regularization, we can also add a formatting penalty to the model that encourages following the format. Echoing the previous discussion on the reward choice, we highlight here that in general checking if the model complies with the formatting requirement is much simpler than checking if the answers are matched.
@@ 7 Related work
Chain-of-thought as latent variable modeling.
@@
Despite the appeal of a full ELBO formulation, it is rarely implemented in practice due to the requirement of learning the posterior distribution. Indeed, despite the formulation of Hoffman et al. (2024) they ended up approximating the posterior with MCMC, which effectively made use of an explicit reward to filter samples. This also marks a key difference from our work - we do not apply any explicit reward scoring throughout our algorithmic design and practical implementation. In addition, Hu et al. (2024) has proposed a more systemic hierarchical latent variable modeling view of chain-of-thought. Similar to our motive, Sordoni et al. (2024) optimized an ELBO inspired objective for prompt selection, where they did not resort to an external reward.
Evidence lower bound and RL.
The connections between evidence lower bound and RL has been extensively studied in both the variational inference (Ranganath et al., 2016; Blei et al., 2017) and RL community (Levine, 2018; O’Donoghue et al., 2020). In the RL literature, much of the variational inference view has often been harnessed to better interpret existing algorithms with much focus on the goal-conditional problems, where a single reward is assigned at the end of a trajectory, akin to the RLHF case (Andrychowicz et al., 2017; Eysenbach et al., 2020; Tang and Kucukelbir, 2021). Our formulation also naturally incorporates the tighter multi-sample lower bound (Burda et al., 2015; Rainforth et al., 2018) as special cases, which has seen little adoption in prior RL literature.
@@
8 Experiments
We empirically validate the relative strength of the lower bound methods against online RL algorithm. We focus on the mathematical reasoning dataset MATH (Hendrycks et al., 2021) where the prompt $x$ asks the model a mathematical question with a short form answer $a^{*}$. We mainly focus on the two algorithmic variants proposed in this work: the lower bound algorithm defined through the gradient estimate in Eqn (5) as well as its multi-sample variant Eqn (7). As a strong baseline, we consider the online policy gradient RL algorithm which applies Sympy (Meurer et al., 2017) to automatically match the answers. The RL algorithm applies leave-one-out for variance reduction, as is commonly practiced (Ahmadian et al., 2024a; Shao et al., 2024). Our main experiments are based on the 8B and 70B model from the Llama 3 model family (Dubey et al., 2024). All algorithmic variants apply identical hyper-parameters such as learning rate, and that they all apply $n=4$ samples for gradient estimations, which we detail in Appendix B.
We highlight again that the RL baseline is at an advantage in this setting, since the reward is fairly accurate and is itself being used as evaluation signals too (Yue et al., 2024). We do not compare with other baselines developed in prior work (e.g., (Hoffman et al., 2024)) as they can be interpreted as variants of online RL algorithms with certain low-level implementation differences.
8.1 Comparion on MATH
During RL training, we use a reward of $r=1$ when there is an answer match and $r=0$ otherwise. Note that lower bound algorithms does not require access to such a reward, but we monitor the reward scores during training. Figure 2 left plot shows the training performance of all baselines. For the x-axis, we use the KL divergence $\mathbb{KL}(\pi_{\theta},\pi_{\text{ref}})$ calculated over the training set. Following the practice in (Gao et al., 2023), we adopt the KL divergence as a certain measure of the optimization budget that the algorithm has consumed. Note that here all experiments are run with the same regularization coefficient $\beta=10^{-3}$ since it achieves a good trade-off for all algorithmic variants over all.
Training performance.
Figure 2 left plot shows that online RL achieves a good KL-performance trade-off on the training set. This is probably not a big surprise since online RL optimizes for the very same objective that we monitor here. In the meantime, lower bound algorithms enjoy reasonable performance: as the policy deviates from the reference policy, the reward performance improves despite not explicitly training for it (in theory lower bound algorithms optimize for a hard string match rather than Sympy match). (2) the multi-sample lower bound algorithm obtains noticeably better performance than the one-sample lower bound baseline, despite using the same $n=4$ generations per update. We will ablate on the impact of parameter $n$ on the performance.
Evaluation.
Figure 2 middle plot shows the evaluation performance on an held-out test set. We note that the reward on the training set is higher than the test set, because the model has been SFT’ed on on the training prompts. For evaluation, observe that the multi-sample lower bound method obtains similar performance as online RL, despite being outperformed during training. We conjecture that this is because online RL tends to overfit the training prompts more significantly, producing a high training reward that does not transfer as well to the evaluation time. This shows that even without training on the reward signal explicitly, lower bound algorithms can obtain a similar evaluation performance as online RL.
Statistical correlation between objectives.

### v2

@@ 5.2 Reinforcement learning
5:     (iii) Update $\theta$ with $n$-sample gradient estimate Eqn (5) or its multi-sample variant Eqn (8).
@@
Algorithm 1 JEPO: Jensen’s evidence lower bound policy optimization (for both single-sample and multi-sample lower bounds)
Assuming access to $n$ i.i.d. trajectories $(y_{i})_{i=1}^{n}\sim\pi_{\theta}(\cdot|x)$, we start with the classic RL policy gradient with leave-one-out baseline (for example, RLOO (Ahmadian et al., 2024))
@@
The proof is provided in Appendix D. Putting $g_{\text{var-reduced-pg}}$ from Eqn (11) and the gradient estimate of the Jensen’s lower bound (Eqn (5)) side-by-side, we identify intriguing similarities. Both gradient estimates employ two terms that update either the chain-of-thought component $\pi_{\theta}(\cdot|x)$ or the answer component $\pi_{\theta}(\cdot|x,c)$, with the only subtle difference being the extra log-transform needed for obtaining the Jensen lower bound. This alludes to the fact that the lower bound gradient has intrinsic built-in variance reduction.
@@ 5.3 Optimizing Jensen’s lower bound with regularization is optimizing a special ELBO
Applying the same derivation as the regular ELBO, the log likelihood $\log p_{\pi_{\theta},\pi_{\text{ref}}}(a|x)$ is lower bounded as
@@
Note that the aggregate objective Eqn (14) can also be optimized via stochastic gradient ascent with standard estimates. We just need to add an additional term associated with the KL regularization, to the original gradient estimate to $\mathcal{L}_{\theta}(x,a^{*})$ defined in Eqn (5). An example of such a gradient estimate is the following
@@
5.4 Practical trade-offs comparing JEPO vs. RL
@@
As discussed earlier, JEPO does not require an external verifiable reward, instead, it can be understood as applying the indicator reward $r(x,y)=\mathds{1}_{\{a=a^{\ast}\}}$. In practice, this can be instantiated as a strict string match float(answer == gt_answer). However, such a reward function will likely induce false negatives, as semantically equivalent generations might be vastly different strings. In practice, a more lenient match is typically applied to remove more false negatives. For example, for math problems (Hendrycks et al., 2021; Yue et al., 2024), usually programmtic checks are implemented to check for equivalence of two short expressions, such that e.g., pi and 3.1415926 might be considered equivalent.
@@
More formally, consider a general reward function $r(x,y)=\text{match}(a,a^{*})$ calculated as a binary match between $a$ and $a^{*}$. We can rewrite the RL objective as $\mathbb{E}[\text{match}(a,a^{*})]$. In order to adapt JEPO to the new RL ojbective, we need to work with the equivalent set $\mathcal{A}\coloneqq\{a|\text{match}(a,a^{*})=1\}$ as well as quantities such as the set probability $\pi_{\theta}(\mathcal{A}|x,c)\coloneqq\sum_{a\in\mathcal{A}}\pi_{\theta}(a|x,c)$. Note that this probability reduces to $\pi_{\theta}(a^{*}|x,c)$ in case we use exact match. In general, computing such probabilities is expensive since we need to enumerate all $a\in\mathcal{A}$ if inverting the match function is feasible at all. As a result, it is challenging to adapt JEPO to generic match function or reward function.
In summary, when a good verifiable reward is available (Sympy vs. string for certain math datasets, see semi-verifiable experiments in Section 9), online RL is at an advantage. There are also cases where good rewards are not readily available and JEPO is a reasonable algorithm. An example is where the ground truth answer takes a rather long form, e.g., see unverifiable experiments in Section 10.
6 Implementation details
We explain the implementation details of the JEPO algorithm in this section. We highlight a few key technical elements for the practical implementation, which we have found to be important in getting the best performance.
Formatting penalty
We find it useful to have an additional RL loss with the reward as $r_{\text{format}}(x,y)=-p$ if $y$ does not follow the formatting requirement (that the identifier phrase the final answer is is in $y$) and zero otherwise. We find that this generally helps stabilize the training process. This is especially useful for small models (8B), where, under temperature sampling, it can often not follow instructions strictly. For large models (70B), we also found that its formatting might be inconsistent after multi-epoch training. We find a value of $p=10$ suffices while smaller values tend to make the training less stable due to weaker penalties.
Per-sequence log probs
For the log-ave-exp operation that defines the lower bound in Eqn (4), it is important to apply the per-sequence log probs without averaging over the sequence length. Concretely, the bound is calculated as
$\log\left(\frac{1}{n}\sum_{j=1}^{n}\sum_{t<|a^{\ast}|}\pi_{\theta}(a_{t}^{\ast}|x,c_{j},a_{<t}^{\ast})\right),$
where $|a^{*}|$ denotes the sequence length of the ground truth $a^{*}$. It is important not to average the sequence level log probs $\log\sum_{t<|a^{\ast}|}\pi_{\theta}(a_{t}^{\ast}|x,c_{j},a_{<t}^{\ast})$ with a length factor of $1/|a^{\ast}|$ as suggested in other contexts (Grinsztajn et al., 2024; Shao et al., 2024). The reason is that the algorithm seeks to make $a^{*}$ more likely, and the sequence level log probs comply with this goal. The length normalization can modify the objective landscape significantly especially when $|a^{*}|$ is large. For example, JEPO algorithm does not work well on the proof data when length normalization is applied.
Advantage normalization
Both the baseline RL and JEPO apply advantage post-processing, following common practice in prior work (Schulman et al., 2017; Dhariwal et al., 2017). For example, in the multi-sample JEPO, the raw advantage for the $i$-th generation is
$A_{i}=\log\left(\frac{1}{n}\sum_{j=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{j})\right)-\tilde{v}_{i},$
where $\tilde{v}_{i}$ is the control variate. A further normalization is applied to the advantage $\tilde{A}_{i}=\text{clip}(A_{i}/\text{std}\left(A\right),-1,1)$ such that the outcome $\tilde{A}_{i}$ is applied in the actual update. Advantage normalization is especially important for JEPO because its raw advantage takes a wider range of values, compared to RL with binary reward.
Weighted supervised learning loss
We also introduce a weighting coefficient for the supervised loss $\beta_{\text{sup}}\geq 0$, which we found useful for ablations. We observe that small values $0\sim 10^{-2}$ tends to work for short-answer applications (e.g., MATH) while a large value $\sim 1$ is important for semi long-form data (e.g., numina and numina-proof), in order to place more weight on the supervised learning loss.
KL-regularization
In early investigations, we found it useful to have a KL regularization at a very small coefficient $\beta=10^{-3}$. The regularization helps prevent formatting collapse, and also prevents the policy from drifting too much in case the updates are noisy (Ziegler et al., 2019; Ouyang et al., 2022).
Figure 3: Verifiable data experiments with MATH. We compare three baselines: online RL with access to the oracle Sympy-based reward and JEPO. In the left plot, we monitor the reward on the training dataset. Online RL obtains the best training time trade-off, followed by multi-sample lower bound and the single-sample lower bound; In the middle plot, we monitor the evaluation on a test set during training. Multi-sample lower bound and online RL obtains similar performance; In the right plot, we graph training reward against the lower bound objectives, averaged over training tokens. The two signals bear positive correlations overall and multi-sample lower bound yields better correlations.
Put together, given $n$ samples, the JEPO update is
$\frac{1}{n}\sum_{i=1}^{n}\left(\left(\tilde{A}_{i}+\tilde{A}_{i}^{\text{(ref)}}\right)\nabla_{\theta}\log\pi_{\theta}(c_{i}|x)\right)$
$+\beta_{\text{sup}}\nabla_{\theta}\log\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{i})\right)$
$-\beta\nabla_{\theta}\mathbb{KL}\left(\pi_{\theta}(\cdot|x),\pi_{\text{ref}}(\cdot|x)\right),$
where $\tilde{A}_{i}^{\text{ref}}$ is the normalized advantage for the formatting penalty. The advantage normalization and weighting coefficient $\beta_{\text{sup}}$ make it such that the ultimate update optimizes for a weighted lower bound with resemblance to to $\beta$-VAE (Higgins et al., 2017). We encourage readers to reference against the online RL implementation in Appendix B to understand its similarities with the JEPO algorithm.
The JEPO loss is only applied to generations with correct format, otherwise, the loss is masked out. The formatting advantage update is applied to all generations. Also, we find that the sequence level normalization with a factor of $1/\left(|c_{i}|+|a^{*}|\right)$ or $1/|c_{i}|$ does not make a significant difference in performance (Shao et al., 2024; Liu et al., 2025).
@@ 7 Related work
Training with unverifiable data
A natural way to generalize RL training to unverifiable data is to make use of LLM feedback, e.g., LLM-as-judge uses LLM to assess the quality of the generated response (Lee et al., 2023; Guo et al., 2024; Yuan et al., 2025). However, despite its conceptual simplicity, LLM-as-judge might not produce reliable assessment for domain-specific or long-form data (Lightman et al., 2023; Petrov et al., 2025). When optimizing against judge scores, it is also more likely to over-optimize (Gao et al., 2023). As a result, in this work we apply LLM-as-judge only for short-form evaluations and not for training.
Closely related to our work is the concurrent VR-CLI (verifiable reward with completion likelihood improvement) (Gurung and Lapata, 2025) where they apply log probs of golden generations as reward. Using our terminology, their approach resembles the first part of the gradient in Eqn (6) of the Jensen’s evidence lower bound. Without a SFT-like component, their update does not optimize for the marginal likelihood only partially. JEPO also applies the multi-sample technique to tighten the lower bound, achieving better empirical performance, which we will demonstrate in Section 8.
Likelihood-based scoring
Prior work showcased the utility of Likelihood-based scoring in filtering of chain-of-thought (Zelikman et al., 2024; Ruan et al., 2025). The algorithms mostly proceed in an iterative fashion akin to expectation-maximization (Moon, 1996), which in theory can also maximize the evidence of the desirable final answers. Complementary to such work, since we extend the training process to fully online RL settings, we forgo the need of variational posteriors which allows for training on unverifiable data at scale. We also demonstrate performance gains beyond short-form answers, which were the main focus of prior work.
Chain-of-thought as latent variable modeling
@@
Despite the appeal of the full ELBO formulation, it is rarely implemented in practice due to the requirement of learning the posterior distribution. Indeed, despite the formulation of Hoffman et al. (2024) they ended up approximating the posterior with MCMC, which effectively made use of an explicit reward to filter samples. This also marks a key difference from our work - we do not apply any explicit reward scoring throughout our algorithmic design and practical implementation. In addition, Hu et al. (2024) has proposed a more systemic hierarchical latent variable modeling view of chain-of-thought. Similar to our motive, Sordoni et al. (2024) optimized an ELBO inspired objective for prompt selection, where they did not resort to an external reward.
Evidence lower bound and RL
The connections between evidence lower bound and RL has been extensively studied in both the variational inference (Ranganath et al., 2016; Blei et al., 2017) and RL community (Levine, 2018; O’Donoghue et al., 2020). In the RL literature, much of the variational inference view has been used to better interpret and improve existing algorithms with much focus on the goal-conditional problems, where a single reward is assigned at the end of a trajectory. Such a setting is quite akin to the RLHF case, where a sequence terminates with a single reward (Andrychowicz et al., 2017; Eysenbach et al., 2020; Tang and Kucukelbir, 2021). Our formulation also naturally incorporates the tighter multi-sample lower bound (Burda et al., 2015; Rainforth et al., 2018) as special cases, which has seen little adoption in prior RL literature.
8 Experiments with verifiable data
We start by comparing JEPO against RL baselines on verifiable data. We focus on the mathematical reasoning dataset MATH (Hendrycks et al., 2021) where the prompt $x$ asks the model a mathematical question with a short form answer $a^{*}$. We study the two algorithmic variants proposed in this work: the JEPO defined through the gradient estimate in Eqn (5) as well as its multi-sample variant Eqn (8). As a strong baseline, we consider the online policy gradient RL algorithm which applies Sympy (Meurer et al., 2017) to automatically match the answers. The RL algorithm applies leave-one-out for variance reduction, as is commonly practiced (Ahmadian et al., 2024; Shao et al., 2024). Our main experiments are based on the 8B and 70B model from the Llama 3.1 model family (Dubey et al., 2024). All algorithmic variants apply identical hyper-parameters such as learning rate, and that they all apply $n=4$ samples for gradient estimations, which we detail in Appendix B.
The RL baseline is at an advantage in this setting, since the reward is of high quality and is itself being used as evaluation signals too (Yue et al., 2024). We do not compare with other baselines developed in prior work (e.g., (Hoffman et al., 2024)) as they can be interpreted as variants of online RL algorithms with relatively minor algorithmic differences.
@@ 8 Experiments with verifiable data
8.1 Comparison on MATH
During RL training, we use a reward of $r=1$ when there is an answer match and $r=0$ otherwise. Note that JEPO does not require access to such a reward, but we monitor the reward scores during training. Figure 3 left plot shows the training performance of all baselines. For the x-axis, we use the KL divergence $\mathbb{KL}(\pi_{\theta},\pi_{\text{ref}})$ calculated over the training set. Following the practice in (Gao et al., 2023), we adopt the KL divergence as a certain measure of the optimization budget that the algorithm has consumed. Note that here all experiments are run with the same regularization coefficient $\beta=10^{-3}$ since it achieves a good trade-off for all algorithmic variants over all.
Training performance
Figure 3 left plot shows that online RL achieves a good KL-performance trade-off on the training set. This is probably not a big surprise since online RL optimizes for the very same objective that we monitor here. In the meantime, JEPO enjoys reasonable performance: as the policy deviates from the reference policy, the reward performance improves despite not explicitly training for it (in theory JEPO optimizes for a hard string match rather than Sympy match). (2) the multi-sample JEPO obtains noticeably better performance than the one-sample lower bound baseline, despite using the same $n=4$ generations per update. We will ablate on the impact of parameter $n$ on the performance.
Evaluation
Figure 3 middle plot shows the evaluation performance on an held-out test set. We note that the reward on the training set is higher than the test set, because the model has been SFT’ed on on the training prompts. For evaluation, observe that the multi-sample lower bound method obtains similar performance as online RL, despite being outperformed during training. We conjecture that this is because online RL tends to overfit the training prompts more significantly, producing a high training reward that does not transfer as well to the evaluation time. This shows that even without training on the reward signal explicitly, JEPO can obtain a similar evaluation performance as online RL.
Statistical correlation between objectives